# Rustam Ibragimova,b    Jihyun Kimc    Anton Skrobotovd a Imperial College Business School b New Economic School

### v2

@@ 7 Conclusion
Table 9: Size and power for the discrete time models, structural break in volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.6 4.9 5.0 4.5 5.0 5.0 4.3 5.0 5.0
$\tau(\check{\beta}_{e})$ 4.8 4.7 4.8 4.6 4.6 4.8 4.9 5.1 5.4
q=4 3.3 3.1 3.0 2.2 2.5 3.0 2.4 2.7 3.0
$t_{8}(\tau(\check{\beta}_{o}))$ 3.4 3.7 3.5 2.2 2.7 3.0 2.0 2.7 3.2
$t_{12}(\tau(\check{\beta}_{o}))$ 4.2 4.1 3.7 2.2 2.8 3.2 1.8 2.6 3.0
$t_{16}(\tau(\check{\beta}_{o}))$ 4.3 4.4 4.4 2.5 2.7 3.3 1.6 2.3 3.0
IVX 32.7 32.5 31.2 32.1 33.5 34.3 32.4 34.1 35.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 9.8 10.1 11.0 7.9 8.9 9.0 9.0 11.4 12.0
$\tau(\check{\beta}_{e})$ 10.2 10.2 10.4 8.0 8.3 8.8 10.4 11.8 12.4
$t_{8}(\tau(\check{\beta}_{o}))$ 15.1 15.3 15.5 5.5 6.7 7.5 5.8 8.1 9.0
$t_{12}(\tau(\check{\beta}_{o}))$ 21.6 22.5 22.1 7.1 9.0 9.3 6.1 8.9 9.9
$t_{16}(\tau(\check{\beta}_{o}))$ 26.3 28.5 28.5 9.0 10.3 12.1 6.6 9.6 10.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 19.3 20.0 20.0 14.0 14.7 15.8 20.0 24.4 26.2
$t_{8}(\tau(\check{\beta}_{o}))$ 34.2 34.0 34.5 11.9 14.1 15.1 13.3 18.6 19.8
$t_{12}(\tau(\check{\beta}_{o}))$ 48.3 50.4 50.2 18.2 20.2 21.7 16.1 21.9 23.3
$t_{16}(\tau(\check{\beta}_{o}))$ 59.6 61.5 62.9 22.8 26.5 28.3 18.2 24.6 26.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 29.4 29.7 30.0 21.9 24.4 25.8 33.9 40.5 43.3
$t_{8}(\tau(\check{\beta}_{o}))$ 50.5 50.7 50.9 21.5 24.0 25.3 24.4 31.9 33.6
$t_{12}(\tau(\check{\beta}_{o}))$ 69.6 71.3 71.2 33.1 35.7 37.3 30.8 38.3 40.3
$t_{16}(\tau(\check{\beta}_{o}))$ 81.2 83.1 83.2 41.5 46.4 47.7 36.5 44.0 47.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 38.5 38.7 38.8 31.8 35.8 37.5 49.0 58.0 60.4
$t_{8}(\tau(\check{\beta}_{o}))$ 63.2 63.5 63.0 31.9 34.5 36.6 36.1 45.2 47.7
$t_{12}(\tau(\check{\beta}_{o}))$ 82.2 83.8 83.8 49.2 51.5 52.3 46.4 54.8 56.8
$t_{16}(\tau(\check{\beta}_{o}))$ 91.6 92.8 92.7 60.3 64.8 65.0 54.6 62.8 65.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 45.5 45.7 46.3 41.9 47.2 48.4 62.4 71.8 74.2
$t_{8}(\tau(\check{\beta}_{o}))$ 71.5 71.7 70.9 41.7 45.3 46.4 46.5 56.7 59.3
$t_{12}(\tau(\check{\beta}_{o}))$ 89.5 90.3 90.3 62.4 64.9 64.7 60.0 68.1 69.8
$t_{16}(\tau(\check{\beta}_{o}))$ 96.2 96.8 96.6 74.1 77.3 78.2 70.3 76.7 78.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 10: Size and power for the discrete time models, regime switching in volatility, MA(2) innovation
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 2.6 2.9 3.5 2.7 2.5 3.0 3.4 3.4 4.0
$t_{12}(\tau(\check{\beta}_{o}))$ 2.4 2.3 3.2 1.8 2.0 2.5 2.3 2.6 3.2
$t_{16}(\tau(\check{\beta}_{o}))$ 2.3 2.1 2.8 1.7 1.8 2.3 2.1 2.0 2.4
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 10.2 10.8 12.1 11.0 11.5 12.3 16.2 17.4 19.0
$t_{12}(\tau(\check{\beta}_{o}))$ 13.3 13.8 15.0 11.5 10.7 11.8 17.0 17.1 18.3
$t_{16}(\tau(\check{\beta}_{o}))$ 16.1 16.0 18.8 12.9 12.0 12.9 17.7 15.4 18.0
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 26.2 27.6 28.7 28.6 30.4 30.1 41.8 45.1 47.2
$t_{12}(\tau(\check{\beta}_{o}))$ 37.5 39.1 40.4 34.4 33.2 33.4 46.9 47.9 48.9
$t_{16}(\tau(\check{\beta}_{o}))$ 47.6 49.3 50.2 41.3 38.2 38.3 53.6 49.7 50.2
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 45.8 46.7 47.9 49.7 51.9 50.3 65.8 70.3 72.3
$t_{12}(\tau(\check{\beta}_{o}))$ 63.9 65.3 66.4 60.3 58.8 58.2 74.7 75.4 76.2
$t_{16}(\tau(\check{\beta}_{o}))$ 76.6 77.6 78.2 70.9 67.5 65.9 82.3 80.3 78.4
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 62.3 62.9 63.6 66.5 68.0 67.3 80.4 84.3 86.0
$t_{12}(\tau(\check{\beta}_{o}))$ 81.2 82.7 83.6 79.0 78.2 76.9 89.0 90.0 90.2
$t_{16}(\tau(\check{\beta}_{o}))$ 91.1 91.9 92.4 88.4 85.9 84.8 94.6 93.5 93.3
@@
$t_{8}(\tau(\check{\beta}_{o}))$ 73.9 75.2 75.6 78.3 79.5 78.8 88.7 91.8 92.9
$t_{12}(\tau(\check{\beta}_{o}))$ 91.0 91.9 92.2 89.7 89.0 88.2 95.7 95.9 96.1
$t_{16}(\tau(\check{\beta}_{o}))$ 96.9 97.1 97.4 95.7 94.6 94.2 98.3 98.0 97.9
@@
Table 11: Size and power for the discrete time models, regime switching in volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 3.8 3.9 4.3 4.3 5.1 5.3 4.6 4.8 4.9
$\tau(\check{\beta}_{e})$ 4.0 4.2 4.3 4.8 4.9 4.7 4.9 4.8 4.5
$t_{8}(\tau(\check{\beta}_{o}))$ 3.5 3.5 3.5 2.3 2.3 2.5 2.3 2.4 3.2
$t_{12}(\tau(\check{\beta}_{o}))$ 3.1 3.4 3.2 1.8 2.0 2.6 2.0 1.9 2.6
$t_{16}(\tau(\check{\beta}_{o}))$ 3.5 3.2 3.4 1.6 1.8 2.1 1.6 1.8 2.0
IVX 15.6 13.6 13.4 9.8 10.5 11.1 9.6 11.2 12.4
$\tau(\check{\beta}_{e})$ 8.5 8.7 8.7 10.5 11.4 11.9 12.7 14.4 14.9
$t_{8}(\tau(\check{\beta}_{o}))$ 16.0 15.5 17.4 8.2 7.9 8.0 8.8 9.8 11.2
$t_{12}(\tau(\check{\beta}_{o}))$ 20.7 21.9 23.2 9.3 8.9 9.2 9.7 10.4 11.3
$t_{16}(\tau(\check{\beta}_{o}))$ 26.5 26.4 28.5 10.8 9.8 11.1 10.5 10.1 10.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 16.2 16.8 16.9 22.9 24.8 24.0 28.5 34.4 35.7
$t_{8}(\tau(\check{\beta}_{o}))$ 37.7 38.7 40.8 21.0 20.2 19.5 22.9 25.2 28.1
$t_{12}(\tau(\check{\beta}_{o}))$ 53.2 53.8 56.8 26.3 25.9 24.4 28.0 29.0 29.3
$t_{16}(\tau(\check{\beta}_{o}))$ 65.2 66.9 69.0 34.5 31.2 32.7 33.5 31.7 32.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 25.4 25.9 26.3 38.3 41.5 41.5 47.6 59.1 60.6
$t_{8}(\tau(\check{\beta}_{o}))$ 59.4 59.6 61.8 37.5 36.3 34.6 42.3 44.5 47.6
$t_{12}(\tau(\check{\beta}_{o}))$ 78.1 78.7 81.1 49.4 48.4 45.9 52.1 53.0 53.6
$t_{16}(\tau(\check{\beta}_{o}))$ 88.3 89.8 90.7 62.2 58.6 58.4 62.0 59.9 59.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 34.3 35.1 35.0 53.1 58.4 58.5 64.7 77.3 79.3
$t_{8}(\tau(\check{\beta}_{o}))$ 74.1 74.3 76.0 53.0 52.1 50.5 59.3 62.3 65.2
$t_{12}(\tau(\check{\beta}_{o}))$ 90.3 91.1 92.0 69.3 68.4 65.7 72.1 72.5 73.0
$t_{16}(\tau(\check{\beta}_{o}))$ 96.6 97.2 97.6 82.1 80.2 78.8 82.8 80.5 79.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\tau(\check{\beta}_{e})$ 42.3 42.4 42.3 64.5 71.5 71.8 76.0 87.6 89.9
$t_{8}(\tau(\check{\beta}_{o}))$ 83.8 83.7 84.8 66.6 65.6 63.7 71.9 75.0 77.3
$t_{12}(\tau(\check{\beta}_{o}))$ 95.8 96.1 96.7 82.8 81.9 79.9 85.1 85.1 85.0
$t_{16}(\tau(\check{\beta}_{o}))$ 99.0 99.3 99.4 92.7 91.5 90.0 93.0 91.2 90.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 12: Empirical results

### v3

@@ 7 Conclusion
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table 8: Size and Power for Discrete-Time Models with Regime Switching in Volatility and MA(1) Innovations
@@
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits regime switching.
Table 9: Empirical Results on Stock Return Predictability (Two-Sided Tests)